OpenStudy (anonymous):

3 5/7 = ? 38/7 26/7 105/7 35/7

OpenStudy (anonymous):

26/7

OpenStudy (anonymous):

Just do 3*7+5. You got 3 wholes (21/7) and add 5/7 to it which is 26/7.

OpenStudy (anonymous):

I used a calculator

OpenStudy (anonymous):

oh, okay. thanks!

OpenStudy (theviper):

\[3\frac{5}{7}\]\[\implies \frac{26}{7}\]

OpenStudy (anonymous):

Since any number times 1 equals the same number: 3 = 3*1 Now 1 can be written as (1/1), (2/2), (3/3), etc. You want to turn 3 into a fraction with 7 as the denominator, so: \[3*\frac{7}{7} = \frac{21}{7}\] Add that to 5/7

OpenStudy (anonymous):

@Suckatmath94 Please don't just give out the answer. That's against the code of conduct and doesn't help anyone to learn math.